One grade of tea worth $3.20 a pound is to be mixed with another grade worth $2 per pound to make 20 pounds that will sell for $2.72 per pound. How much of each grade of tea must be used?
Thanks a ton!!!

Answer by stanbon(75887)

One grade of tea worth $3.20 a pound is to be mixed with another grade worth $2 per pound to make 20 pounds that will sell for $2.72 per pound. How much of each grade of tea must be used?
----------------------------------------
Equation:
value + value = value
3.20x + 2(20-x) = 2.72*20
Multiply thru by 100 to get:
320x + 200*20 - 200x = 272*20
120x = 72*20
x = 12 lb (amount of $3.20 tea in the mixture)
---------------
20-x = 8 lb (amount of $2.00 tea in the mixture)
=======================================================
Cheers,
Stan H.